# Breaking Information Cocoons: A Hyperbolic Graph-LLM Framework for Exploration and Exploitation in Recommender Systems

### v2

@@ Abstract
Modern recommender systems often create information cocoons, restricting users’ exposure to diverse content. A key challenge lies in balancing content exploration and exploitation while allowing users to adjust their recommendation preferences. Intuitively, this balance can be modeled as a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration. However, existing approaches face two fundamental limitations: Euclidean methods struggle to capture hierarchical structures, while hyperbolic methods, despite their superior hierarchical modeling, lack semantic understanding of user and item profiles and fail to provide a principled mechanism for balancing exploration and exploitation. To address these challenges, we propose HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems. Our framework introduces two key innovations: (1) a hierarchical-aware graph-LLM mechanism that jointly aligns textual descriptions with user-item collaborative information in hyperbolic space, and (2) a hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs. Extensive experiments demonstrate that HERec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, effectively mitigating information cocoons.
@@ 1 INTRODUCTION
As online platforms grow in size dramatically, users are overwhelmed by an influx of information, creating information cocoons that restrict their exposure to diverse and novel content (Kunaver & Požrl, 2017; Castells et al., 2021).
@@
Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data (He et al., 2020; Koren et al., 2009), where users with similar historical preferences are expected to share future interests (Chen et al., 2021). However, these methods face two critical challenges in improving user experience. First, existing models struggle to capture the underlying hierarchical structure inherent in user-item networks (Zheng et al., 2017; Unger & Tuzhilin, 2020), a critical factor for enhancing performance and personalization. Second, these models fail to effectively balance the exploration-exploitation trade-off. For instance, a jazz enthusiast may want to explore rock music, but current models primarily focus on exploitation—recommending only familiar genres—without adequately supporting exploration (as illustrated in Figure 1).
@@
Hyperbolic space has emerged as a promising approach for modeling hierarchical structures and power-law distributions (Krioukov et al., 2010; Yang et al., 2022b; Sun et al., 2021; Yang et al., 2022a), where its exponential volume growth naturally aligns with hierarchical, scale-free structures (Ganea et al., 2018; Chami et al., 2019; Liu et al., 2019; Zhang et al., 2021; Yang et al., 2023). In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries. This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance (Yang et al., 2022a; Sun et al., 2021; Yang et al., 2022b).
@@
Despite these advantages, current hyperbolic recommender systems still face key limitations. First, they struggle with a lack of semantic understanding of user and item profiles. While large language models (LLMs) offer a potential solution, their text encoders operate in Euclidean space, making it challenging to integrate LLM-derived semantic features with hyperbolic collaborative signals. This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively. Second, hyperbolic recommender systems lack a principled approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) (Wilson et al., 2021). This limitation reduces their ability to adapt to the exploration-exploitation trade-off, ultimately diminishing both the utility and diversity of recommendations.
@@
Proposed Methods. To address these challenges, we propose HERec, a hyperbolic graph-LLM framework designed to balance exploration and exploitation in recommender systems. Our approach introduces two key innovations: (1) a hierarchical-aware graph-LLM alignment mechanism, which jointly aligns textual descriptions with user-item collaborative information in hyperbolic space, supported by theoretical gradient analysis demonstrating its advantages, and (2) a novel hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs. Unlike conventional approaches that require manually defined hierarchies, we propose a hyperparameter-free clustering mechanism theoretically optimized by Dasgupta’s cost (Dasgupta, 2016). This approach automatically discovers hierarchical structures without requiring predefined hyperparameters, allowing for a more adaptive and principled organization of user preferences.
@@
To validate the effectiveness of HERec, we conducted extensive experiments across multiple dimensions. Our model consistently outperformed baselines in both utility and diversity metrics, marking a significant achievement as it is the first model to excel in both aspects simultaneously. For a fair comparison, we enhanced Euclidean baselines with feature information, yet our model still demonstrated superior performance, validating the advantage of our designed alignment in hyperbolic space. Additionally, we conducted a detailed analysis on head and tail items, showing that HERec effectively boosts recommendations for tail items, thus enhancing diversity. An ablation study, along with a hierarchical representation structure analysis and decoder module, confirms HERec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
@@
Contributions. Our primary contributions focus on three key aspects: capturing collaborative signals, integrating semantic information, and modeling the underlying hierarchical structure. The proposed graph-LLM alignment framework unifies collaborative and semantic information in hyperbolic space, while a hierarchical representation structure organizes user preferences, allowing for dynamic adjustment between exploration and exploitation. Moreover, through extensive experiments, we demonstrate that HERec achieves state-of-the-art performance across both utility and diversity metrics, making it the first model to excel in both criteria simultaneously. To facilitate reproducibility and further research, we open-source our model implementation at: https://github.com/Martin-qyma/HERec.
Figure 1: Exploration and exploitation in hyperbolic space.
@@ 2 Related Work
Graph Collaborative Filtering. Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role. Unlike traditional collaborative filtering techniques, such as matrix factorization (Koren et al., 2009), GNNs capitalize on the relational structure in user-item interactions, enabling a more nuanced understanding of complex dependencies. Prominent graph-based methods, such as NGCF (Wang et al., 2019), SGL (Wu et al., 2021), and LightGCN (He et al., 2020), have been widely adopted and have consistently achieved state-of-the-art performance. Nevertheless, these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution in user-item networks.
Large Language Models for Recommendation. Large language models (LLMs) are increasingly used in recommender systems for their strong capability in understanding textual information. Some approaches leverage LLMs directly for downstream recommendation tasks (Bao et al., 2023; Li et al., 2023; Lyu et al., 2023; Zheng et al., 2024; Zhang et al., 2023), while others utilize LLMs as an auxiliary tool for collaborative filtering (Ren et al., 2024; Xi et al., 2024; Wei et al., 2024). However, these approaches require substantial inference time and are computationally inefficient. In our HERec, we use language models exclusively to improve data quality, eliminating the need for repeated inference during each recommendation by utilizing hyperbolic embeddings for the final output.
Hyperbolic Representation Learning. To address the limitations of Euclidean space in modeling hierarchical structures, researchers have investigated the use of hyperbolic space (Chami et al., 2019; Liu et al., 2019; Yang et al., 2022b; Zhang et al., 2021), where distances grow exponentially as embeddings move away from the origin. This property makes hyperbolic geometry particularly well-suited for modeling user-item interactions by more accurately capturing latent hierarchies and dependencies. Consequently, hyperbolic representation learning has gained considerable attention in the recommender systems domain (Ying et al., 2018; Vinh Tran et al., 2020; Sun et al., 2021; Yang et al., 2022a), demonstrating improved performance over traditional Euclidean models.
However, none of these hyperbolic models fully leverage the rich semantic information embedded in language-based datasets. While some approaches incorporate knowledge graphs as supplementary information (Tai et al., 2021; Du et al., 2022), these models primarily focus on refining hierarchical structures rather than achieving a deeper semantic understanding of each item or user.
@@ 3 PRELIMINARIES
Embeddings learned from graph collaborative filtering are by default in the Euclidean space, which is unable to capture tree-like structures. To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
Riemannian Manifold. A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold that features an inner product $g_{x}:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$ at each point $x\in\mathcal{M}$. This inner product, known as the Riemannian metric, defines geometric characteristics of the space such as angles and curve lengths. The curvature of Riemannian manifolds results in different types of geometries: elliptic geometry with positive curvature, Euclidean geometry with zero curvature, and hyperbolic geometry with negative curvature. In this work, we focus on hyperbolic geometry, characterized by negative curvature. Hyperbolic space can be described by several equivalent models, each offering unique properties while being mathematically equivalent. The Lorentz model, also referred to as the hyperboloid model, is one of the commonly used representations of hyperbolic space.
Lorentz Model. An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$. Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined
@@
A tangent space is an $n$-dimensional vector space that approximates Euclidean space to hyperbolic space. Specifically, given a point $x\in\mathbb{H}_{\kappa}^{n}$, there exists a tangent space $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ that approximates $\mathbb{H}_{\kappa}^{n}$:
@@
We also need to define transitions between $\mathbb{H}_{\kappa}^{n}$ and $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$, which give rise to the exponential and logarithmic maps. For $x\in\mathbb{H}_{\kappa}^{n}$ and $v\in\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ with $v\neq\mathbf{0}$ and $y\neq x$, there exists a unique geodesic $\gamma:[0,1]\to\mathbb{H}_{\kappa}^{n}$ such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=v$. The exponential map $\exp_{x}:\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}\to\mathbb{H}_{\kappa}^{n}$ is given by $\exp_{x}(v)=\gamma(1)$:
@@
where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ denotes the Lorentzian norm of $v$. The logarithmic map $\log_{x}$ is the inverse of the exponential map $\exp_{x}$, and the distance between $x$ and $y$ in $\mathbb{H}^{n}_{\kappa}$ is:
@@
We choose $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ as the reference point for all the operations, making the transitions reproducible. In this work, we set $\kappa=1$, resulting in a curvature of $-1$. Therefore, in the following discussions, we would omit $\kappa$ for simplicity.
@@ 4.1 Hyperbolic Graph Collaborative Filtering
Similar to Euclidean space, hyperbolic graph collaborative filtering captures user-item dependencies via message passing mechanism. In our method, we use the Lorentz representation for both users and items.
Hyperbolic Embedding Initialization. As discussed in section 3, we fix the origin $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ and use it as the reference point. We adopt the Gaussian sampling method for the initialization of embeddings in Euclidean space, after that, by injecting them into the tangent space of the reference point, we derive the initial hyperbolic embeddings of both users and items:
@@
where $e_{u},e_{i}$ are Gaussian distributed initialized Euclidean embeddings, and $z_{u}^{0},z_{i}^{0}$ are corresponding embeddings in the tangent space of the origin point. $h_{u}^{0},h_{i}^{0}$ therefore stands for the resulting initialized hyperbolic embeddings.
Hyperbolic Message Passing. Message passing between each node is realized through hyperbolic neighborhood aggregation, which is used to extract explicit user-item interaction. Given user $u$, item $i$, as well as their neighborhood $\mathcal{N}_{u},\mathcal{N}_{i}$, their embeddings at layer $l$ in the tangent space are updated as follows:
$z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\ z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$ (7)
where $|\mathcal{N}_{u}|$ and $|\mathcal{N}_{i}|$ represent the number of neighbors of $u$ and $i$, respectively. The final tangent space embedding is derived by summing up embeddings at each layer:
@@
Since $z_{u}$ and $z_{i}$ are on the tangent space, we still need to project it into hyperbolic space to derive the hyperbolic embeddings:
@@
Hyperbolic Margin Ranking and Informative Mechanism. To address the limitations of Euclidean-based approaches, we adopt the hyperbolic margin ranking loss (Sun et al., 2021), an extension of Bayesian Personalized Ranking (BPR) (Rendle et al., 2012). This loss pulls user-item pairs with observed interactions closer while pushing negative pairs apart, preventing the collapse seen in Euclidean space by incorporating a margin:
@@
To further refine this pull-push process, we integrate a geometric-aware mechanism (Yang et al., 2022a) combining Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS). HAML dynamically adjusts the margin based on hyperbolic geometry, assigning a larger margin to pairs near the origin:
@@
where $e_{u,0}$ and $e_{i,0}$ are the zeroth coordinate elements of $e_{u}$ and $e_{i}$, respectively, corresponding to the embeddings norms in the hyperbolic space. HINS enhances negative sampling by prioritizing candidates close to the positive item, iteratively identifying the most informative negative samples. This method leverages hyperbolic geometry to effectively model head and tail items, improving utility performance in recommendations.

### v3

@@ Abstract
Modern recommender systems often create information cocoons, restricting users’ exposure to diverse content. A key challenge lies in balancing content exploration and exploitation while allowing users to adjust their recommendation preferences. Intuitively, this balance can be modeled as a tree-structured representation, where depth search facilitates exploitation and breadth search enables exploration. However, existing approaches face two fundamental limitations: Euclidean methods struggle to capture hierarchical structures, while hyperbolic methods, despite their superior hierarchical modeling, lack semantic understanding of user and item profiles and fail to provide a principled mechanism for balancing exploration and exploitation. To address these challenges, we propose HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems. Our framework introduces two key innovations: (1) a semantic-enhanced hierarchical mechanism that aligns rich textual descriptions processed by large language models (LLMs) with collaborative information directly in hyperbolic space, allowing for more nuanced updates that respect the underlying hierarchical structure in user-item profiles; (2) an automatic hierarchical representation by optimizing Dasgupta’s cost, which discovers hierarchical structures without requiring predefined hyperparameters, enabling user-adjustable exploration-exploitation trade-offs. Extensive experiments demonstrate that HERec consistently outperforms both Euclidean and hyperbolic baselines, achieving up to 5.49% improvement in utility metrics and 11.39% increase in diversity metrics, effectively mitigating information cocoons.
@@ 1 INTRODUCTION
As online platforms grow in size dramatically, users are overwhelmed by an influx of information, creating information cocoons that restrict their exposure to diverse and novel content [16, 4].
@@
Traditional recommendation approaches have primarily relied on collaborative filtering and interaction data [11, 14], where users with similar historical preferences are expected to share future interests [6]. However, these methods face two critical challenges in improving user experience. First, existing models struggle to capture the underlying hierarchical structure inherent in user-item networks [48, 27], a critical factor for enhancing performance and personalization. Second, these models fail to effectively balance the exploration-exploitation trade-off. For instance, a jazz enthusiast may want to explore rock music, but current models primarily focus on exploitation—recommending only familiar genres—without adequately supporting exploration.
@@
Hyperbolic space has emerged as a promising approach for modeling hierarchical structures and power-law distributions [15, 37, 25, 36], where its exponential volume growth naturally aligns with hierarchical, scale-free structures [10, 5, 18, 46, 38]. In hyperbolic space, distances increase exponentially from the origin, allowing popular items or exploratory users to be positioned closer to the origin, while niche items or users with focused interests spread toward the boundaries. This spatial configuration enables hyperbolic graph neural network (GNN) models to perceive hierarchical relationships effectively, enhancing recommendation performance [36, 25, 37].
@@
Despite these advantages, current hyperbolic recommender systems still face key limitations. First, they struggle with a lack of semantic understanding of user and item profiles. While large language models (LLMs) offer a potential solution, their text encoders operate in Euclidean space, making it challenging to integrate LLM-derived semantic features with hyperbolic collaborative signals. This mismatch leaves existing hyperbolic models vulnerable to noise in collaborative information, as they fail to capture semantic insights effectively. Second, hyperbolic recommender systems lack a principled approach to balancing exploration (introducing new content) and exploitation (focusing on existing preferences) [32]. This limitation reduces their ability to adapt to the exploration-exploitation trade-off, ultimately diminishing both the utility and diversity of recommendations.
@@
Proposed Method. To address these challenges, we propose HERec, a hyperbolic graph-llm framework designed to balance exploration and exploitation in recommender systems. Our approach introduces two key innovations: (1) a hierarchical-aware graph-LLM alignment mechanism, which jointly aligns textual descriptions processed by LLMs with user-item collaborative information in hyperbolic space, supported by theoretical gradient analysis demonstrating its adaptiveness; (2) a novel hierarchical representation structure that enables user-adjustable exploration-exploitation trade-offs. Unlike conventional approaches that require manually defined hierarchies, we propose a hyperparameter-free clustering mechanism theoretically optimized by Dasgupta’s cost [7]. This approach automatically discovers hierarchical structures without requiring predefined hyperparameters, allowing for a more adaptive and principled organization of user preferences.
@@
To validate the effectiveness of HERec, we conducted extensive experiments across multiple dimensions. Our model consistently outperformed baselines in both utility and diversity, marking a significant achievement as it is the first model to excel in both aspects simultaneously. For a fair comparison, we also enhanced Euclidean baselines with feature information, yet our model still demonstrated superior performance. Additionally, we conducted a detailed analysis on head and tail items, showing that HERec effectively boosts recommendations for tail items, thus enhancing diversity. An ablation study, along with a hierarchical representation structure analysis module, confirms HERec’s capability to leverage both semantic meaning and collaborative information, providing deeper insights into the advantages of our approach.
@@
Contributions. Our primary contributions focus on three key aspects: capturing collaborative signals, integrating semantic information, and modeling the underlying hierarchical structure. The proposed graph-llm alignment framework unifies collaborative and semantic information in hyperbolic space, while a hierarchical representation structure organizes user preferences, allowing for dynamic adjustment between exploration and exploitation. Moreover, through extensive experiments evaluating both utility and diversity metrics, we demonstrate that HERec achieves state-of-the-art performance across utility and diversity metrics, making it the first model to excel in both criteria simultaneously. We have made our model implementation available at: https://github.com/Martin-qyma/HERec.
@@ 2 Related Work
Graph Collaborative Filtering. Graph Collaborative Filtering has emerged as a dominant approach in recommender systems, with graph neural networks (GNNs) playing a central role. Unlike traditional collaborative filtering techniques, such as matrix factorization [14], GNNs capitalize on the relational structure in user-item interactions, enabling a more nuanced understanding of complex dependencies. Prominent graph-based methods, such as NGCF [30], SGL [33], and LightGCN [11], have been widely adopted and have consistently achieved state-of-the-art performance. Nevertheless, these models are constrained to Euclidean space, which may fall short in capturing the underlying power-law distribution in user-item networks.
Large Language Models for Recommendation. Large language models (LLMs) are increasingly used in recommender systems for their strong capability in understanding textual information. Some approaches leverage LLMs directly for downstream recommendation tasks [1, 17, 19, 47, 45], while others utilize LLMs as an auxiliary tool for collaborative filtering [23, 34, 31]. However, these approaches require substantial inference time and are computationally inefficient. In our HERec, we use language models exclusively to improve data quality, eliminating the need for repeated inference during each recommendation by utilizing hyperbolic embeddings for the final output.
Hyperbolic Representation Learning. To address the limitations of Euclidean space in modeling hierarchical structures, researchers have investigated the use of hyperbolic space [5, 18, 37, 46], where distances grow exponentially as embeddings move away from the origin. This property makes hyperbolic geometry particularly well-suited for modeling user-item interactions by more accurately capturing latent hierarchies and dependencies. Consequently, hyperbolic representation learning has gained considerable attention in the recommender systems domain [41, 29, 25, 36], demonstrating improved performance over traditional Euclidean models.
However, none of these hyperbolic models fully leverage the rich semantic information embedded in language-based datasets. While some approaches incorporate knowledge graphs as supplementary information [26, 9], these models primarily focus on refining hierarchical structures rather than achieving a deeper semantic understanding of each item or user.
@@ 3 PRELIMINARIES
Embeddings learned from graph collaborative filtering are by default in the Euclidean space, which is unable to capture hierarchical structures. To address the problem, we introduce how to derive these embeddings in a hyperbolic space.
Riemannian Manifold. A Riemannian manifold $(\mathcal{M},g)$ is a smooth manifold with an inner product $g_{x}$ at each point $x\in\mathcal{M}$, defining geometric properties such as angles and curve lengths. Its curvature determines the geometry: positive (elliptic), zero (Euclidean), or negative (hyperbolic). We focus on hyperbolic geometry, modeled using the Lorentz (hyperboloid) representation.
Lorentz Model. An $n$-dimensional Lorentz manifold with negative curvature $-1/\kappa$ $(\kappa>0)$ is described as the Riemannian manifold $(\mathbb{H}^{n}_{\kappa},g_{\mathcal{L}})$. Here, $\mathbb{H}^{n}_{\kappa}=\{x\in\mathbb{R}^{n+1}:\langle x,x\rangle_{\mathcal{L}}=-\kappa,x_{0}>0\}$, $g_{\mathcal{L}}=\eta$ with $\eta=\mathbf{I}_{n}$ except $\eta_{0,0}=-1$, and $\langle\cdot,\cdot\rangle_{\mathcal{L}}$ represents the Lorentzian inner product, which is defined as:
@@
A tangent space is an $n$-dimensional vector space that locally approximates hyperbolic space. For a point $x\in\mathbb{H}_{\kappa}^{n}$, the tangent space $\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}$ is defined as:
@@
This space is orthogonal to $x$ under the Lorentzian inner product, providing accurate local approximation. Transitions between hyperbolic space and tangent space are facilitated by exponential and logarithmic maps, essential for gradient-based optimization. The exponential map $\exp_{x}:\mathcal{T}_{x}\mathbb{H}_{\kappa}^{n}\to\mathbb{H}_{\kappa}^{n}$ projects a tangent vector onto hyperbolic space along a geodesic:
@@
where $\|v\|_{\mathcal{L}}=\sqrt{\langle v,v\rangle_{\mathcal{L}}}$ is the Lorentzian norm. The logarithmic map $\log_{x}$ is the inverse of $\exp_{x}$ and retrieves the tangent vector for a point $y\in\mathbb{H}_{\kappa}^{n}$:
@@
We use $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ as the reference point. For simplicity, we set $\kappa=1$, giving a curvature of $-1$, and omit $\kappa$ in subsequent discussions.
@@ 4.1 Hyperbolic Graph Collaborative Filtering
Similar to Euclidean space, hyperbolic graph collaborative filtering captures user-item dependencies via message passing mechanism. We apply the Lorentz representation for both users and items.
Hyperbolic embedding initialization. As discussed in section 3, we fix the origin $\mathbf{o}:=\{\sqrt{\kappa},0,\ldots,0\}\in\mathbb{H}^{n}_{\kappa}$ and use it as the reference point. Embeddings are initialized in Euclidean space using Gaussian sampling and projected onto the tangent space of $\mathbf{o}$ to derive their hyperbolic representations:
@@
where $e_{u},e_{i}$ are Gaussian-initialized Euclidean embeddings, and $z_{u}^{0},z_{i}^{0}$ are their tangent space projections. $h_{u}^{0},h_{i}^{0}$ denote the resulting hyperbolic embeddings.
Hyperbolic Message Passing. User-item interactions are captured via hyperbolic neighborhood aggregation. For user $u$, item $i$, and their neighborhoods $\mathcal{N}_{u},\mathcal{N}_{i}$, tangent space embeddings at layer $l$ are updated as:
$z_{u}^{l}=z_{u}^{l-1}+\sum_{i\in\mathcal{N}_{u}}\frac{1}{|\mathcal{N}_{u}|}z_{i}^{l-1},\quad z_{i}^{l}=z_{i}^{l-1}+\sum_{u\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}z_{u}^{l-1}$ (7)
where $|\mathcal{N}_{u}|$ and $|\mathcal{N}_{i}|$ are the neighborhood sizes. The final tangent embeddings are obtained by summing over all layers:
@@
Since $z_{u}$ and $z_{i}$ are on the tangent space, we still need to project them into hyperbolic space to derive the hyperbolic embeddings:
@@
Hyperbolic Margin Ranking and Informative Mechanism. To address the limitations of Euclidean-based approaches, we adopt the hyperbolic margin ranking loss [25], an extension of Bayesian Personalized Ranking (BPR) [24]. This loss pulls user-item pairs with observed interactions closer while pushing negative pairs apart, preventing the collapse seen in Euclidean space by incorporating a margin:
@@
To further refine this pull-push process, we integrate a geometric-aware mechanism [36] combining Hyperbolic-aware Margin Learning (HAML) and Hyperbolic Informative Negative Sampling (HINS). HAML dynamically adjusts the margin based on hyperbolic geometry, assigning a larger margin to pairs near the origin:
@@
where $e_{u,0}$ and $e_{i,0}$ are the zeroth coordinates of $e_{u}$ and $e_{i}$, representing their norms. HINS enhances negative sampling by prioritizing candidates near the positive item, leveraging hyperbolic geometry to better model head and tail items, ultimately improving recommendation performance.